# Sub-shot noise sensitivity via deformed four-headed kitten states

### v1

@@ III.1 Deformed Coherent States
Figure 5: PND for SA and AS cases are represented by $P_{\ket{\oplus}}(n)$ and $P_{\ket{\ominus}}(n)$, respectively. (a) $p=0$, $q=0$; (b) $p=4$, $q=2$; (c) $p=4$, $q=4$; (d) $p=0$, $q=2$; (e) $p=4$, $q=2$; and (f) $p=4$, $q=4$. We use $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$ in all cases, and distributions are normalized to unity.
Figure 6: The overlap function $|F_{\square}(c_{0})|^{2}$ is represented as a blue dashed curve and $|F_{\blacksquare}(c_{0})|^{2}$ as a red solid line, with $p=4$ and $q=4$. The horizontal black dashed line depicts the case where $p=0$ and $q=0$; consequently, the overlap is simply between two compass states.
III.2 Photon Statistics
@@
$P_{\ket{\oplus}}(n)=\frac{N_{\oplus}[(q+n)!]^{2}\kappa}{n![(q+n-p)!]^{2}}~{}\text{and}~{}P_{\ket{\ominus}}(n)=\frac{N_{\ominus}n!\kappa}{[(n-p)!]^{2}}$ (18)
@@
$\kappa:=|\alpha|^{2(q-p+n)}\text{e}^{-|\alpha|^{2}}.$ (19)
@@
Let us now analyze these distributions. Here, Figs. 5(a)-(c) show the PND for the SA case of the coherent state for different situations based on the varying amount of added and subtracted photons. As depicted in Fig. 5(a), PND with $p=0$ and $q=0$ exhibits a Poissonian distribution, which obviously corresponds to a coherent state. In Fig. 5(b), when $p=4$ photons are added and set $q=0$, the Poissonian distribution shifts to higher values of $n$, with the peak now occurring at a larger mean photon number. In Fig. 5(c), the Poissonian distribution shifts to lower values of $n$, as with an increment in the subtracting photons to $q=4$, and the number of added photons is kept at $p=4$ as in the prior instance. These cases reflects that in the SA examples, a higher $p$ corresponds to a higher average photon number, whereas a higher $q$ corresponds to a lower average photon number in the resultant state.
@@
The PND for AS case is shown in Figs. 5(d)-(f) with a few $p$ and $q$ values. As illustrated in Fig. 5(d), for $q=2$ and $p=0$, the statistics of PND stay constant, showing that raising the number of subtracted photons has no influence on the average photon number of the states when applied directly to a coherent state. As observed in Fig. 5(e), increasing the number of added photons while maintaining the number of subtracted photons constant, that is, with $p=4$ and $q=2$, drives the Poissonian distribution to a larger $n$. In Fig. 5(f), the number of subtracted photons is increased to $q=4$ without changing $p$, and it is observed that subtracting photons from a coherent state has no influence on the related PND. This is an additional confirmation of how photon subtraction operations keeps PND of this case invariant. Hence, photon addition in both the SA and AS situations increases the average photon number in the subsequent states; however, the AS case maintains the distribution at higher mean photon values, implying that this situation has higher average photon numbers than the SA case.
In the upcoming sections, we present and explore the SA and AS cases of the compass states. We provide a comprehensive analysis that builds on the foundational concepts introduced above. This detailed examination presents the discussion about the characteristics and implications of these specific cases, offering insights into their behavior and significance within the framework of our study.
@@
Figure 7: (a)-(c) show the Wigner function of the SA case of the compass state, and (d)-(f) show the analogous AS cases. (a) $p=12$, $q=12$, (b) $p=24$, $q=12$, (c) $p=24$, $q=20$, (d) $p=12$, $q=12$, (e) $p=24$, $q=12$, and (f) $p=24$, $q=20$. In all cases, we set $c_{0}=1$.
III.3 Variants of a Kitten State
@@
In §III.1, we noticed that the SA and AS variants of a standard coherent state exhibit fascinating phase-space properties that surpass those of the original coherent states. It is observed that the sequence in which photon operations are applied significantly affects the outcome, resulting in two distinct quantum states with different phase-space characteristics for the SA and AS cases. We now extend these concepts to other quantum states of interest in the sections that follow.
@@
We now introduce SA and AS cases of our interest. For our SA scenario, $p$ photons are added to the compass state $\ket{\diamondsuit}$ as given in Eq. (5), followed by the subtraction of $q$ photons. In the AS scenario, $q$ photons are first subtracted from the compass state, and then $p$ photons are added. These photon operations on the compass state produce new quantum states, which may display distinct phase-space characteristics.
@@
|$\ket{\square}:=N_{\square}^{-\nicefrac{{1}}{{2}}}\hat{a}^{q}\hat{a}^{\dagger p}\sum^{4}_{i=1}\ket{\alpha_{i}}$ (20)
@@
$N_{\square}=$ $(-1)^{p+q}\sum^{4}_{i,j=1}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{p-n,q}\left[\text{i}\alpha_{j},\text{i}\alpha^{*}_{i}\right]$
×$\times H_{p-n,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right]$ (21)
@@
|$\ket{\blacksquare}:=N_{\blacksquare}^{-\nicefrac{{1}}{{2}}}\hat{a}^{\dagger p}\hat{a}^{q}\sum^{4}_{i=1}\ket{\alpha_{i}},$ (22)
@@
$N_{\blacksquare}=$ $\sum^{4}_{i,j=1}(-1)^{n}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{p-n,q}\left[\alpha_{j},\alpha^{*}_{i}\right]$
×$\times H_{p-n,q}\left[\alpha^{*}_{i},\alpha_{j}\right]$ (23)
represents the normalization factor for this case.
We now compare the original compass state with the proposed variants by assessing their overlap. This overlap indicates how distinct the proposed states are from the original compass state. By optimizing the parameters to maximize this distinction, we will analyze the characteristics of the states with these optimized parameters.
Let us evaluate the overlap between our SA case and the compass state. The overlap is represented by $F_{\square}(c_{0}):=\braket{\square}{\diamondsuit}$ and is calculated as
$F_{\square}(c_{0})=(-\text{i})^{p+q}\left[N_{\square}N_{\diamondsuit}\right]^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\square_{i}}\bra{\square_{j}}}$ (24)
@@
$F_{\ket{\square_{i}}\bra{\square_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{p,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right],$ (25)
@@
$F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{p,q}\left[\alpha^{*}_{i},\alpha_{j}\right],$ (27)
@@
In Fig. 6, the black dashed line represents the overlap between two same compass states, and recalling that the parameter $c_{0}$ measures the mutual separation between the coherent states, as discussed in the §II.2, this parameter has a strong impact on the characteristics of the original compass state, that is, the existence of the sub-Planck structures and enhanced sensitivity is associated with higher $c_{0}$ values. For example, the scenario represented in Fig. 1(a) for comparatively smaller $c_{0}$ does not have the capacity to exhibit these traits, and this particular case is named as four-headed kitten state.
Note that in our plots, we only investigated the overlaps for the $p=q=4$ case, because it is understood here that for other higher values of $p$ and $q$, our proposed states will hold more difference different than the compass state for smaller $c_{0}$ values, as depicted here; thus, this result is meaningful for other settings of $p$ and $q$. In these plots, it is clear that greater $c_{0}$ values cause the blue dashed and red solid lines to coincide with the horizontal dashed black line, suggesting that the respective SA and AS cases revived to the original compass state. However, when parameter $c_{0}$ is smaller, our proposed SA and AS cases diverge significantly from the original compass state, as illustrated in the plot, where the blue dashed and red solid curves are much lower than the black horizontal dashed line.
In the following, we assume $c_{0}=1$ and then investigate the detailed phase-space characteristics of the associated four-headed cat state using varied photon addition and subtraction operations. Here, we simply refer to these consequent states as SA and AS cases of this kitten state, and our analysis is divided into the subsections follows. In §III.4, we demonstrate and examine the phase-space characteristics of these quantum states. In §III.5, we analyze the sensitivity to phase-space displacement for each example.
@@
Figure 8: The central phase-space structures of the cases depicted in Fig. 7. (a) $p=12$, $q=12$, (b) $p=24$, $q=12$, (c) $p=24$, $q=20$, (d) $p=12$, $q=12$, (e) $p=24$, $q=12$, and (f) $p=24$, $q=20$. In all cases, we set $c_{0}=1$.
III.4 Analysis of Distinctive Phase-Space Features
@@
To this point, we have introduced the SA and AS cases of our interest, each producing distinct quantum states. In these instances, photon addition and subtraction processes appeared to have a main impact on their characteristics, and in this section, we particularly employ the Wigner function to investigate the corresponding phase space of these quantum states. To obtain the Wigner function for each case, Eq. (1) is employed, and we denote $W_{\ket{\square}}(\beta)$ and $W_{\ket{\blacksquare}}(\beta)$ as the corresponding Wigner functions of SA and AS cases of the kitten state, respectively.
@@
Let us now examine the Wigner distributions for each scenario. For the SA case of our kitten state introduced in Eq. (20), the Wigner distribution is calculated as follows:
@@
$W_{\ket{\square_{i}}\bra{\square_{j}}}(\beta):=$ $W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta)\sum^{p}_{n=0}\Gamma H_{p-n,q}\left[\text{i}\Omega^{*}_{j},\text{i}\alpha_{i}\right]$
×$\times H_{p-n,q}\left[-\text{i}\Omega_{i},-\text{i}\alpha^{*}_{j}\right]$ (29)
@@
$W_{\ket{\blacksquare}}(\beta)=\frac{1}{N_{\blacksquare}}\sum^{4}_{i,j=1}W_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\beta),$ (31)
@@
$W_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\beta):=$ $W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta)\sum^{p}_{n=0}\Gamma H_{p-n,q}\left[\Omega^{*}_{j},\alpha_{i}\right]$
×$\times H_{p-n,q}\left[\Omega_{i},\alpha^{*}_{j}\right].$ (32)
@@
The relevant Wigner functions are presented in Fig. 7, with Fig. 8 illustrating the central phase-space features of each case, where for both figures the cases labelled with (a)-(c) exhibit the SA, while (d)-(f) provides equivalent AS cases of the kitten state. It is readily apparent that our SA and AS instances achieve substantially distinct phase-space characteristics. Interestingly, a central sub-Planck structure is identified in each of the cases outlined. We focus on the significance of the sub-Planck structure in these states and examine this particular phase-space feature in detail.
@@
Photons addition and subtraction appeared to have a significant effects in the Wigner distribution of the corresponding states. Note that the the original compass states, represented in Fig. 1(a) with $c_{0}=1$, do not exhibit a sub-Planck structure. As we stated before, this particular situation is also known as a kitten state. This implies that appearance of the sub-Planck structures in our cases is attributed to the photon operations involved. Specifically, it is observed that as the parameter $p$, which represents the number of added photons, increases, the size of sub-Planck structure reduces uniformly in both the SA and AS cases. This is evident by comparing the scenarios presented in Figs. 8(a) and 8(b) for the SA case, and then Figs. 8(d) and 8(e) is for corresponding AS cases, where increasing $p$ with constant $q$, clearly depicts this impact. We have already noticed in §III.1 that photons subtraction operation applied to a coherent state has no effects in the phase space, but when combined with the photons addition operation, this operation has also the significant effects, that is, an increase in the parameter $q$ denoting the number of photons subtractions causes to reduce the size of the sub-Planck structure, as observed in Figs. 8(b) and 8(c) for the SA case and then Figs. 8(e) and 8(f) are representing this impact for the AS cases.
A notable change in the sub-Planck structures is observed when transitioning from the Zurek compass state to the SA and AS cases, that is, some of these cases display nearly isotropic sub-Planck structures resemble to those of the compass state, whereas others, especially as shown in Figs. 8(e) and 8(f), are quite isotropic. The isotropic behavior of these cases is clearly evident, as demonstrated by the circular structure around the origin. This indicates that, for these scenarios, the sub-Planck structure is uniformly constrained in all directions of the phase space. Note that our variants for $q=0$ are simply photon-added kitten states, as it is understood that such states may also exhibit similar phase-space features, whereas the case when $p=0$ is irrelevant to the current discussion because the ineffectiveness of the parameter $q$ without parameter $p$ is obvious from the discussion. Furthermore, the SA and AS cases with differing average photon numbers, with the AS example having a higher average photon number, appear to exhibit distinct sub-Planck structures, demonstrated by the associated phase space intensity figures.
In summary, photon addition and subtraction in a smaller compass state, also referred to as a four-component (or four-headed) kitten state, result in a significantly larger compasslike state that maintains a sub-Planck structure in its phase space. Interestingly, this transition from the compass state to our states has the advantage of making the ensuing sub-Planck structures isotropic in some situations. Our variants may also exhibit comparable metrological capacity as the compass state, which we will explore in the next section by analyzing their sensitivity to phase-space displacement.
@@
Figure 9: Overlap between a state and its slightly translated version with (a)-(c) depict SA examples, whereas (d)-(f) exhibit comparable AS cases: (a) $p=12$, $q=12$, (b) $p=24$, $q=12$, (c) $p=24$, $q=20$, (d) $p=12$, $q=12$, (e) $p=24$, $q=12$, and (f) $p=24$, $q=20$. In all situations, we set $c_{0}=1$ and normalize the intensity plots to unity.
III.5 Enhancement in Sensitivity

### v2

@@ III.1 Deformed coherent states
Figure 6: PND for SA and AS cases are represented by $P_{\ket{\oplus}}(n)$ and $P_{\ket{\ominus}}(n)$, respectively. (a) $r=0$, $q=0$; (b) $r=4$, $q=2$; (c) $r=4$, $q=4$; (d) $r=0$, $q=2$; (e) $r=4$, $q=2$; and (f) $r=4$, $q=4$. We use $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$ in all cases, and distributions are normalized to unity.
III.2 Fluctuations in photon numbers
@@
$P_{\ket{\oplus}}(n)=\frac{N_{\oplus}[(q+n)!]^{2}\kappa}{n![(q+n-r)!]^{2}}~{}\text{and}~{}P_{\ket{\ominus}}(n)=\frac{N_{\ominus}n!\kappa}{[(n-r)!]^{2}}$ (19)
@@
$\kappa:=|\alpha|^{2(q-r+n)}\text{e}^{-|\alpha|^{2}}.$ (20)
@@
Let us now analyze these distributions. Here, Figs. 6(a)-(c) show the PND for the SA case of the coherent state for different situations based on the varying amount of added and subtracted photons. As depicted in Fig. 6(a), PND with $r=0$ and $q=0$ exhibits a Poissonian distribution, which obviously corresponds to a coherent state. In Fig. 6(b), when $r=4$ photons are added and set $q=0$, the Poissonian distribution shifts to higher values of $n$, with the peak now occurring at a larger mean photon number. In Fig. 6(c), the Poissonian distribution shifts to lower values of $n$, as with an increment in the subtracting photons to $q=4$, and the number of added photons is kept at $r=4$ as in the prior instance. These cases reflects that in the SA examples, a higher $r$ corresponds to a higher average photon number, whereas a higher $q$ corresponds to a lower average photon number in the resultant state.
Figure 7: The overlap function $|F_{\square}(c_{0})|^{2}$ is represented as a blue dashed curve and $|F_{\blacksquare}(c_{0})|^{2}$ as a red solid line, with $r=4$ and $q=4$. The horizontal black dashed line depicts the case where $r=0$ and $q=0$; consequently, the overlap is simply between two compass states.
@@
The PND for AS case is shown in Figs. 6(d)-(f) with a few $r$ and $q$ values. As illustrated in Fig. 6(d), for $q=2$ and $r=0$, the statistics of PND stay constant, showing that raising the number of subtracted photons has no influence on the average photon number of the states when applied directly to a coherent state. As observed in Fig. 6(e), increasing the number of added photons while maintaining the number of subtracted photons constant, that is, with $r=4$ and $q=2$, drives the Poissonian distribution to a larger $n$. In Fig. 6(f), the number of subtracted photons is increased to $q=4$ without changing $r$, and it is observed that subtracting photons from a coherent state has no influence on the related PND. This is an additional confirmation of how photon subtraction operations keeps PND of this case invariant. Hence, photon addition in both the SA and AS situations increases the average photon number in the subsequent states; however, the AS case maintains the distribution at higher mean photon values, implying that this situation has higher average photon numbers than the SA case.
@@
Figure 8: (a)-(c) show the Wigner function of the SA case of the four-headed kitten state, and (d)-(f) show the analogous AS cases. (a) $r=12$, $q=12$, (b) $r=24$, $q=12$, (c) $r=24$, $q=20$, (d) $r=12$, $q=12$, (e) $r=24$, $q=12$, and (f) $r=24$, $q=20$. In all cases, $c_{0}=1$.
III.3 Impact of multiphoton operations
@@
In §III.1, we noticed that the SA and AS variants of a standard coherent state exhibit fascinating phase-space properties that surpass those of the original coherent states. It is observed that the sequence in which photon operations are applied significantly affects the outcome, resulting in two distinct quantum states with different phase-space characteristics. We now extended this basic notion to the main quantum states of the present work.
@@
First, for our SA scenario, $r$ photons are added to the kitten state $\ket{\diamondsuit}$ as given in Eq. (5), followed by the subtraction of $q$ photons. In the AS scenario, $q$ photons are first subtracted from the compass state, and then $r$ photons are added. These photon operations on the kitten state produce new quantum states with distinct phase-space characteristics.
@@
|$\ket{\square}:=N_{\square}^{-\nicefrac{{1}}{{2}}}\hat{a}^{q}\hat{a}^{\dagger r}\sum^{4}_{i=1}\ket{\alpha_{i}}$ (21)
@@
$N_{\square}=$ $(-1)^{r+q}\sum^{4}_{i,j=1}\sum^{r}_{n=0}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{r-n,q}\left[\text{i}\alpha_{j},\text{i}\alpha^{*}_{i}\right]$
×$\times H_{r-n,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right]$ (22)
@@
|$\ket{\blacksquare}:=N_{\blacksquare}^{-\nicefrac{{1}}{{2}}}\hat{a}^{\dagger r}\hat{a}^{q}\sum^{4}_{i=1}\ket{\alpha_{i}},$ (23)
@@
$N_{\blacksquare}=$ $\sum^{4}_{i,j=1}\sum^{r}_{n=0}(-1)^{n}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha_{i}\alpha_{j})^{q}$
$(\alpha_{j})^{r-n}(\alpha^{*}_{i})^{r-n}$ (24)
represents the normalization factor for this case. When photon subtraction is combined with a photon addition operator, the effect of the parameter $q$ is highly apparent, as we will show later in the following discussion.
We now compare the original compass state with the proposed variants by assessing their overlap. This overlap indicates how distinct the proposed states are from the original compass state. The overlap between our SA case and the compass state is represented by $F_{\square}(c_{0}):=\braket{\square}{\diamondsuit}$ and is calculated as
$F_{\square}(c_{0})=(-\text{i})^{r+q}\left[N_{\square}N_{\diamondsuit}\right]^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\square_{i}}\bra{\square_{j}}}$ (25)
@@
$F_{\ket{\square_{i}}\bra{\square_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{r,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right],$ (26)
@@
$F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha^{*}_{i})^{q}(\alpha_{j})^{r},$ (28)
@@
In Fig. 7, the black dotted line represents the overlap between two same compass states, and recalling that the parameter $c_{0}$ measures the distance from the origin of the phase space, as discussed in §II.2, this parameter has a strong impact on the characteristics of the original state, that is, the existence of the sub-Planck structures, and enhanced sensitivity is associated with higher $c_{0}$ values. For example, the scenario represented in Fig. 1(a) for comparatively smaller values of $c_{0}$ does not have the capacity to exhibit these traits, and this particular case is named as a four-headed kitten state in the present work.
Note that in our plots, we only investigate the overlaps for the certain situation when $r=q=4$, because it is understood here that for other higher values of $r$ and $q$, our proposed states keep the difference than the compass state provided that $c_{0}$ is smaller, as depicted here; simply, this result is meaningful for other settings of $r$ and $q$. In these plots, it is clear that greater $c_{0}$ values push the blue dotted and red solid lines closer to the horizontal black dotted line, implying that the SA and AS cases revived to the original compass state. In our work, we set $c_{0}=1$ and investigate the corresponding phase-space characteristics of the four-headed cat state using multiple options of photon addition and subtraction operations. Here, we simply refer to these consequent states as SA and AS cases of this kitten state. Here, note that our AS case is the eigenstate of the operator $\hat{a}^{4}$, i.e., $\hat{a}^{4}\ket{\blacksquare}=\ket{\diamondsuit}$. This shows that choosing the parameter $q$ as a multiple of 4 ($q=4k$ with $k\in\mathbb{N}^{+}$) corresponds to a photon-added version of the compass state. When $p=0$, this reduces to the compass state $\ket{\diamondsuit}$ itself.
@@
Figure 9: Zoom in on the central phase-space structures of the cases shown in Fig. 8. The cases on the left are SA, whereas the cases on the right are AS. (a) $r=12$, $q=12$; (b) $r=24$, $q=12$; (c) $r=24$, $q=20$; (d) $r=12$, $q=12$; (e) $r=24$, $q=12$ (f) $r=24$, $q=20$. In each example, $c_{0}=1$.
Figure 10: The zoom in on the center phase-space structures of the AS case of the four-headed kitten state with multiple parameter $r$ and $q$ values. (a) $r=1$, $q=1$; (b) $r=5$, $q=5$; (c) $r=9$, $q=9$; (d) $r=16$, $q=10$; (e) $r=22$, $q=10$; (f) $r=22$, $q=18$. For all cases, $c_{0}=1$.
III.4 Unique phase-space features
@@
In the previous section, we have introduced the SA and AS cases of our interest, each producing distinct quantum states. In these instances, photon addition and subtraction processes appeared to have a main impact on their phase-space characteristics, and in this section, we particularly employ the Wigner function to investigate the corresponding phase space of these quantum states. To obtain the Wigner function for each case, Eq. (1) is employed, and we denote $W_{\ket{\square}}(\beta)$ and $W_{\ket{\blacksquare}}(\beta)$ as the corresponding Wigner functions of SA and AS cases of the kitten state, respectively.
@@
Let us now examine the Wigner distributions for each scenario. For the SA case of our four-headed kitten state introduced in Eq. (21), the Wigner distribution is calculated as follows:
@@
$W_{\ket{\square_{i}}\bra{\square_{j}}}(\beta):=$ $W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta)\sum^{r}_{n=0}\Gamma H_{r-n,q}\left[\text{i}\Omega^{*}_{j},\text{i}\alpha_{i}\right]$
×$\times H_{r-n,q}\left[-\text{i}\Omega_{i},-\text{i}\alpha^{*}_{j}\right]$ (30)
@@
$W_{\ket{\blacksquare}}(\beta)=\frac{(-1)^{r}}{N_{\blacksquare}}\sum^{4}_{i,j=1}W_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\beta),$ (32)
@@
$W_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}(\beta):=$ $(\alpha_{i}\alpha^{*}_{j})^{q}W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta)\sum^{r}_{n=0}(-1)^{n}\Gamma(2\beta-\alpha_{i})^{r-n}$
×$\times(\alpha_{j}-2\beta^{*})^{r-n}.$ (33)
@@
The relevant Wigner functions are presented in Figs. 8, with Fig. 9 illustrating the zoom in of the central phase-space features of each case, where for both figures the cases labelled with (a)-(c) exhibit the SA, while (d)-(f) provides equivalent AS cases of the kitten state. With specific selections of the parameters $r$ and $q$, Fig. 10 offers a further illustration of the Wigner function of corresponding AS scenarios.
It is readily apparent that our SA and AS instances achieve substantially distinct phase-space characteristics as compared to the four-headed kitten state. Interestingly, a central sub-Planck structure is identified in each of the cases outlined. We focus on the significance of the sub-Planck structure in these states and examine this particular phase-space feature in detail.
@@
Photon addition and subtraction appeared to have significant effects on the Wigner distribution of the corresponding states. Note that the original four-headed kitten state, represented in Fig. 1(a) with $c_{0}=1$, does not exhibit a sub-Planck structure. This implies that the appearance of the sub-Planck structures in our cases is attributed to the photon operations involved. Specifically, it is observed that as the parameter $r$, which represents the number of added photons, increases, the size of the sub-Planck structure reduces uniformly in both the SA and AS cases. This is evident by comparing the scenarios presented in Figs. 9(a) and 9(b) for the SA case, and then Figs. 9(d) and 9(e) are for corresponding AS cases, where increasing $r$ with constant $q$, clearly depicts this impact. Furthermore, increasing the number of photon subtractions $q$ increases the size of the sub-Planck structure in the SA case, as observed by comparing cases depicted in Figs. 8(b) and 8(c), while it has no effect in the AS case, as shown in Figs. 8(e) and 8(f). Note that all the examples of AS described in Figs. 8 and 9 correspond to the merely photon-added case of the compass state, as for these situations $q$ are the multiple of 4, where the corresponding four-headed kitten state is the eigenstate of the operator $\hat{a}^{4}$.
We find that photon addition generally reduces the size of sub-Planck structures in both SA and AS scenarios, while photon subtraction increases their size in the SA case but has no effect on their size in the AS case. To further illustrate this for the AS case, we examine situations where the number of photon subtractions is not a multiple of 4, as shown in Fig. 10. Increasing the photon addition to the state as previously noted also reduces the size of the sub-Planck structures for this specific illustration (Figs. 10(d) and 10(e)), while as expected, varying the photon subtraction in these cases has no impact on the size of the sub-Planck structures (Figs. 10(e) and 10(f)), but in some specific instances, photon subtraction slightly enhances the isotropy of the sub-Planck structures. For example, an enhancement in the isotropy of the central sub-Planck structures of these cases is clearly evident in Figs. 10(a) to 10(c), as demonstrated by the circular structure around the origin. This indicates that, for these scenarios, the sub-Planck structure is uniformly constrained in all directions of the phase space.
@@
Figure 11: Overlap between a state and its slightly translated version with (a)-(c) depict SA examples, whereas (d)-(i) exhibit comparable AS cases: (a) $r=12$, $q=12$, (b) $r=24$, $q=12$, (c) $r=24$, $q=20$, (d) $r=12$, $q=12$, (e) $r=24$, $q=12$, (f) $r=24$, $q=20$, (g) $r=q=1$, (h) $r=q=5$, and (i) $r=q=9$. In all situations, we set $c_{0}=1$ and normalize the intensity plots to unity.
IV Enhancement in sensitivity